# RAG-Reward: Optimizing RAG with Reward Modeling and RLHF

### v1

@@ Abstract
Retrieval-augmented generation (RAG) enhances Large Language Models (LLMs) with relevant and up-to-date knowledge, improving their ability to answer knowledge-intensive questions. It has been shown to enhance both generation quality and trustworthiness. While numerous works have focused on improving retrieval, generation, and evaluation, the role of reward models in reinforcement learning for optimizing RAG and establishing automated benchmarking pipelines remains underexplored. In this paper, we introduce RAG-Reward, a dataset designed to enable hallucination-free, comprehensive, reliable, and efficient RAG. We define four key metrics for assessing generation quality and develop an automated annotation pipeline that leverages multiple LLMs to generate outputs across diverse RAG scenarios. GPT-4o is used to evaluate and construct preference data. Using RAG-Reward, we train reward models and apply reinforcement learning with human feedback (RLHF) to improve LLMs’ effectiveness in RAG. Experimental results show that our reward model achieves state-of-the-art performance on a held-out test set, demonstrating both the effectiveness of our approach and the quality of our dataset. Furthermore, the improved generation quality of the trained policy model highlights the feasibility of using RLHF to enhance RAG pipelines [^1] Our dataset is available at: https://huggingface.co/datasets/HanningZhang/RAG-Reward-Modeling.
@@ 1 Introduction
Figure 1: An overview of our data labeling method and our experiments based on the preference data in RAG Scenario. We use GPT-4o as the judge to evaluate the quality of the generation from multiple models. We then train the reward models and use them for Reinforcement Learning.
@@
However, many open-source LLMs are not yet fully optimized for use in RAG scenarios. Simply applying RAG to these models often leads to suboptimal outcomes. For example, even with access to external knowledge, these LLMs may still generate misinformation (Niu et al., 2024), which can be especially problematic given the timeliness of the generated content. While many LLMs excel in conversational tasks, they frequently struggle in information-intensive scenarios, where both efficiency and simplicity are paramount. Therefore, a solution that can provide effective supervision and evaluation of RAG systems is needed.
@@
Fortunately, significant progress has been made in evaluating the quality of LLM-generated content these days. Unlike many previous benchmark data sets that are created to assess the overall quality of generations (Yang et al., 2018; Kwiatkowski et al., 2019), recent studies suggest that the construction of evaluation data sets tailored to specific domains or scenarios with detailed criteria (Zhu et al., 2024; Friel et al., 2024) can lead to improved in-domain performance. We argue that it is also feasible to build such a preference dataset that reflects the quality of RAG based on multiple key metrics carefully selected and defined by human experts.
@@
At the same time, reward modeling has recently emerged as a widely adopted strategy to align text generation with human values by learning preference signals from human-annotated high-quality data (Ouyang et al., 2022; Lambert et al., 2024). Multiple well-crafted datasets have demonstrated the value of reward models and high-quality preference datasets. For example, HH-RLHF (Bai et al., 2022a) evaluates the helpfulness and harmlessness of the language, PRM800K (Lightman et al., 2023) assesses the stepwise correctness during mathematical reasoning, and RAG-Truth (Niu et al., 2024) detects hallucination occurrence in RAG systems.
@@
While human annotation has proven to be an effective method to evaluate the quality of generated responses and constructing datasets, it is both expensive and time-consuming. These limitations have led many researchers to explore reliable annotation strategies that leverage the capabilities of LLMs. Previous studies have demonstrated the feasibility of using LLMs to assess response quality and build robust reward models (Tan et al., 2024). For example, Bai et al. (2022b) introduced Constitution AI, which uses LLMs to guide the generation of fine-tuning data and provide preference pairs for reinforcement learning. Moreover, LLMs can also be guided by human-defined criteria and examples to provide preference scores tailored to more specific requirements (Sun et al., 2023, 2024).
@@
The use of LLMs for annotation has further proven effective when applied to a broader range of tasks and larger datasets. For example, Cui et al. (2024) introduced Ultra-Feedback, a framework that leverages GPT-4 to assign scores based on attributes such as helpfulness, truthfulness, honesty, and instruction-following ability. LLMs have also shown their effectiveness in providing feedback in a variety of reasoning tasks, including coding and solving mathematical problems (Weyssow et al., 2024; Yuan et al., 2024).
Building on the inspiration from these works, we further harness the capabilities of LLMs to construct a preference dataset tailored to specific RAG frameworks across diverse domains. To achieve this, we develop a novel automatic LLM annotation pipeline, culminating in a high-quality preference dataset, RAG-Reward, explicitly designed for the general RAG task. Figure 1 shows an overview of our framework. Several works are related to this, such as measuring the relevance of query and passage using reward models (Nguyen et al., 2024), and improve trustworthiness of RAG using DPO (Song et al., 2024). However, to the best of our knowledge, we are the first to integrate RAG with the standard reward modeling and RLHF pipeline.
@@
Specifically, we select datasets from domains of Question-Answering, Data-to-Text, and Summarization. First, we sample diverse responses from a pool of 12 open-source and proprietary LLMs, including GPT-4 and the Llama-3 series. For each prompt in the datasets, we randomly select 2 LLMs to generate responses. Next, we use GPT-4o as the judge to compare them based on four key metrics carefully selected by the human experts: hallucination, comprehensiveness, verbosity, and attribution. This enables us to construct preference pairs, consisting of a chosen response and a rejected one, based on majority voting across 4 metrics. Overall, we collect 35K high-quality training samples for the reward model. Evaluations demonstrate that our reward model achieves over 80% accuracy on the held-out test set. Additionally, we develop a policy model using the RAFT algorithm (Dong et al., 2023), leading to notable performance improvements. Our key contributions are summarized as follows:
@@ 2.1 Reward Modeling for Alignment and Reinforcement Learning
Reinforcement Learning from human feedback (RLHF) is a widely used strategy to enhance policy models after the reward model is developed (Kaufmann et al., 2024). RLHF plays a critical role in aligning LLMs with human values and achieving improved performance (Christiano et al., 2023). Proximal Policy Optimization (PPO) is a commonly used algorithm for alignment tasks to enhance the policy models (Schulman et al., 2017), although it is computationally intensive. Consequently, several alternative, more efficient algorithms have been proposed, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024), and Rejection Sampling Fine-tuning (RAFT) (Dong et al., 2023). These approaches have also been widely adopted in state-of-the-art models like Llama-3 (Grattafiori et al., 2024) and Qwen-2 (Yang et al., 2024a).
@@ 2.2 Large Language Models and Retrieval Augmented Generation
In this project, we are the first to systematically construct RAG-scenario datasets and develop reward models, paving the way for evaluating and enhancing the generation quality of LLMs within the RAG framework.
Figure 2: An illustration of our data annotation method. Given a sample and two responses, we prompt GPT-4o to provide a judgment based on each metric separately. We then aggregate the results and construct the pairs.
@@ 3 Dataset Construction
We construct our dataset based on existing RAG datasets to ensure its relevance and applicability. To reflect the diverse use cases of RAG scenarios, we include three common types: Question Answering, Data-to-Text and Summarization. Specifically, we use WebGLM (Liu et al., 2023), Yelp (Yelp, 2021), and XSum (Narayan et al., 2018) as experimental datasets, each dataset corresponding to one of the three RAG scenarios. Examples of these data sets are presented in Table 2.
Figure 3: The percentage contribution of different metrics for the chosen responses. Each bar shows the percentage of the chosen responses are preferred on that metric.
For the WebGLM dataset, LLMs are tasked with reasoning over web-retrieved reference data to answer real-world questions, generating concise responses in a few sentences. For the Yelp dataset, our experiments focus on data from the restaurant category, represented in JSON format. Each sample includes information such as a restaurant’s location and ambiance. Based on the structured JSON input, LLMs generate descriptive text about the restaurant. The XSum dataset contains diverse articles from the British Broadcasting Corporation (BBC), with models tasked with summarizing these articles. These three datasets cover a broad range of circumstances, ensuring that the reward model trained on them can significantly improve the development and evaluation of RAG systems. Table 1 presents the number of data samples used in our experiments.
@@
While the response should be detailed and comprehensive, it should also be concise and straight to the point. Striking the right balance between detail and brevity is essential for delivering informative responses without overwhelming the user.
@@ 3.2 Dataset Labeling
We use GPT-4o (OpenAI et al., 2024) to label the data. An illustration of our labeling methods is shown in Figure 2. Given a question and a pair of responses from different models, we prompt GPT-4o to compare and select the preferred response.
Specifically, we ask GPT-4o to compare the responses based on the four metrics outlined earlier, assessing them individually. The responses are then ranked according to these evaluations, with the response exhibiting more preferred metrics being chosen as the selected response, and the other marked as rejected. In cases where both responses receive an equal number of preferred metrics, the data is discarded. This approach acknowledges real-world scenarios where responses are rarely perfect, and trade-offs are often necessary.
3.3 Dataset Composition
We analyze the composition of the constructed dataset by examining how each metric contributes to the formation of the preference data. Since our selection process is based on the majority of preferred metrics, the chosen responses may not excel across all four dimensions. For example, a chosen response could properly refer to the context, cover all the details specified in the question without hallucinations, but be too verbose. To better understand how each metric influences the overall preference decision, we evaluate the quality of each metric in relation to the chosen responses.
We calculate the preferred percentage of each metric, and the results are shown in Figure 3. Firstly, we observe a high preferred percentage of each metric. Overall, the average preferred rate across the four dimensions is nearly 90%, demonstrating the high quality of the chosen response on the 4 metrics with respect to the evaluated metrics. We also observe varying preferred rates between the metrics. Specifically, hallucination and attribution metrics exhibit extremely high preferred percentages, while comprehensiveness and verbosity show a relatively lower rate. This difference suggests that the reward models trained on our preference data are better able to differentiate issues related to hallucinations and attribution in RAG scenarios. In contrast, the reward model might not prioritize comprehensiveness and verbosity metrics in their evaluations.
@@ 4.1 Self Evaluation
Consist. 92.8 92.6 94.8 93.4
Table 3: The consistency rate for the self-evaluation of the constructed dataset. We measure whether the GPT-4o agrees with the comparison results previously made.
Figure 4: An illustration of our self-evaluation method. For a sample of the constructed data pairs, we provide GPT-4o with the previous judgment, and ask it to re-evaluate the comparison result.
In this subsection, we evaluate the consistency of the evaluations provided by GPT-4o, a key metric commonly used to demonstrate the reliability of the annotations. To objectively assess the quality of our constructed dataset, we design a self-evaluation method that measures the consistency of GPT-4o’s responses. An illustration of this method is shown in Figure 4. Specifically, for each prompt, a chosen response, and a rejected response in the dataset, we prompt GPT-4o to revisit its previous comparison result and verify whether it maintains its original judgment. We define the consistency rate as the proportion of samples where the evaluation results remain unchanged across both assessments. In this experiment, we randomly select 1,000 samples for re-evaluation.
The results of this experiment are presented in Table 3. We observe a very high consistency rate across the three tasks and the overall consistency rate exceeds 90%. The results demonstrate that GPT-4 produces stable and consistent labels for most of the data according to well-defined criteria, and reflects the quality of the dataset.
@@ 4.2 Human Evaluation
WebGLM Yelp XSum Avg
Consist. 0.68 0.65 0.79 0.71
Table 4: The consistency rate between human evaluation and the GPT-4o labeled dataset.
@@
internlm2-7b-reward (Cai et al., 2024) 77.2 83.0 74.0 78.1
UltraRM-13b (Cui et al., 2024) 78.8 82.0 73.2 78.0
FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2024) 74.4 82.8 75.2 77.5
Eurus-RM-7b (Yuan et al., 2024) 75.4 82.6 74.0 77.3
llama-3-tulu-2-8b-uf-mean-rm (Ivison et al., 2024) 75.0 78.2 72.2 75.1
Llama-3-OffsetBias-RM-8B (Park et al., 2024) 68.6 79.0 72.0 73.2
URM-LLaMa-3.1-8B (Lou et al., 2024) 66.4 79.0 67.6 71.0
GRM-Llama3-8B-rewardmodel-ft (Yang et al., 2024b) 66.0 74.0 72.8 70.9
QRM-Llama3.1-8B-v2 (Dorka, 2024) 69.2 72.2 69.2 70.2
Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) 67.4 72.8 67.8 69.3
GRM-Llama3.2-3B-rewardmodel-ft (Yang et al., 2024b) 67.4 68.0 70.6 68.7
Table 5: The evaluation results of the existing reward models on the 3 tasks. They achieve SOTA performance on chatting, safety, and reasoning evaluation, but fall short on RAG tasks.
We also performed human evaluations to assess the alignment of GPT annotations with human preferences. Specifically, we randomly select 100 samples with paired responses from each dataset and ask the annotators to evaluate using the same pipeline described in Section 3.2 and illustrated in Figure 2. Annotators compare the responses based on each metric and determine the preference pair. We calculate the agreement ratio between the human annotators and GPT-4o in the preference pairs, and the results are shown in Table 4. We observe a promising overall agreement rate of 71%, showing the potential of AI-assisted labeling on multiple RAG tasks.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enhances Large Language Models (LLMs) with relevant and up-to-date knowledge, improving their ability to answer knowledge-intensive questions. It has been shown to enhance both generation quality and trustworthiness. While numerous works have focused on improving retrieval, generation, and evaluation, the role of reward models in reinforcement learning for optimizing RAG remains underexplored. In this paper, we introduce RAG-Reward, a framework designed to develop reward models to enable hallucination-free, comprehensive, reliable, and efficient RAG. We define four key metrics to assess generation quality and develop an automated benchmarking pipeline to evaluate the outputs of multiple LLMs across a variety of RAG scenarios. Using RAG-Reward, we train reward models and apply reinforcement learning with human feedback (RLHF) to improve LLMs’ effectiveness in RAG. Experimental results demonstrate that our reward model achieves state-of-the-art performance in automatic benchmarking and aligns closely with human evaluations. Furthermore, the improved generation quality of the trained policy model highlights the feasibility and efficiency of using RLHF to enhance RAG outputs  [^1] Our dataset is available at: https://huggingface.co/datasets/HanningZhang/RAG-Reward-Modeling .
@@ 1 Introduction
Figure 1: An overview of our data labeling method and our experiments based on the preference data in RAG Scenario. We use o3-mini as the judge to evaluate the quality of the generation from multiple models. We then train the reward models and use them for Reinforcement Learning.
@@
However, many open-source LLMs are not yet fully optimized for use in RAG scenarios. Simply applying RAG to these models often leads to suboptimal outcomes. For example, even with access to external knowledge, these LLMs may still generate misinformation (Niu et al., 2024), which can be especially problematic given the timeliness of the generated content. While many LLMs excel in conversational tasks, they often struggle in information-intensive scenarios, where both trustworthiness and efficiency are paramount Wang et al. (2024). Therefore, a solution that can provide effective supervision and evaluation of RAG systems is needed.
@@
Fortunately, significant progress has been made in evaluating the quality of LLM-generated content these days. Unlike many previous benchmark data sets that are created to assess the overall quality of generations (Yang et al., 2018; Kwiatkowski et al., 2019), recent studies suggest that the construction of evaluation data sets tailored to specific domains or scenarios with detailed criteria (Zhu et al., 2024; Friel et al., 2024) can lead to improved in-domain performance. (Friel et al., 2025) shows the feasibility to build such a preference dataset that reflects the RAG quality based on multiple key metrics carefully selected and defined by human experts.
@@
At the same time, reward modeling has recently emerged as a widely adopted strategy to align text generation with human values by learning preference signals from human-annotated high-quality data (Ouyang et al., 2022; Lambert et al., 2024). Multiple well-crafted datasets have demonstrated the value of reward models and high-quality preference datasets. For example, HH-RLHF (Bai et al., 2022a) evaluates the helpfulness and harmlessness of the language, PRM800K (Lightman et al., 2023) assesses the stepwise correctness during mathematical reasoning. And (Song et al., 2024a) develop hallucination detection models using RAG-Truth (Niu et al., 2024) dataset, and achieves hallucination reduction in RAG output.
@@
While human annotation has proven to be effective in evaluating the quality of the generated responses and constructing datasets, it is both expensive and time-consuming. These challenges have led researchers to explore reliable annotation strategies that leverage the capabilities of LLMs. Previous studies have demonstrated the feasibility of using LLMs to assess response quality and build robust reward models (Tan et al., 2024). For instance, Bai et al. (2022b) introduced Constitution AI, which uses LLMs to guide the generation of fine-tuning data and provide preference pairs for reinforcement learning.
@@
Additionally, LLMs can be guided by human-defined criteria and examples to generate preference scores tailored to specific requirements(Sun et al., 2023, 2024). For example, Cui et al. (2024) introduced Ultra-Feedback, a framework that leverages GPT-4 to assign scores based on attributes such as helpfulness, truthfulness, honesty, and instruction-following ability. LLMs have also shown their effectiveness in providing feedback across various reasoning tasks, including coding and solving mathematical problems (Weyssow et al., 2024; Yuan et al., 2024).
Inspired by these works, we developed RAG-Reward, a framework designed to create reward models that enable hallucination-free, comprehensive, reliable, and efficient RAG. The framework leverages the capabilities of LLMs to evaluate RAG responses across diverse domains, train reward models, and enhance RAG response quality through RLHF. Figure 1 provides an overview of our framework.
Several related works have explored various aspects of this approach, such as using reward models to measure the relevance of queries and passages (Nguyen et al., 2024), improving trustworthiness of RAG using DPO (Song et al., 2024b), leveraging LLM to build RAG benchmark dataset (Friel et al., 2025), and comparing efficiency of various LLMs serving as RAG reward models (Jin et al., 2024). However, to the best of our knowledge, this is the first attempt to optimize RAG with reward modeling and RLHF pipeline.
@@
Specifically, we select datasets from domains of Question-Answering, Data-to-Text, and Summarization. First, we sample diverse responses from a pool of 12 open-source and proprietary LLMs, including GPT-4 and the Llama-3 series. For each prompt in the datasets, we randomly select 2 LLMs to generate responses. Next, we use o3-mini as the judge to compare them based on four key metrics carefully selected by the human experts: hallucination, comprehensiveness, verbosity, and attribution. This enables us to construct preference pairs, consisting of a chosen response and a rejected one, based on the weighted selection across 4 metrics. Overall, we collect 35K high-quality training samples for the reward model training. Evaluations demonstrate that our reward model achieves over 80% accuracy on the held-out test set. Additionally, we develop a policy model using the RAFT algorithm (Dong et al., 2023), leading to notable performance improvements. Our key contributions are summarized as follows:
@@ 2.1 Reward Modeling for Alignment and Reinforcement Learning
Reinforcement Learning from human feedback (RLHF) is a widely used strategy to enhance policy models after the reward model is developed (Kaufmann et al., 2024). RLHF plays a critical role in aligning LLMs with human values and achieving improved performance (Christiano et al., 2023). Proximal Policy Optimization (PPO) is a commonly used algorithm for alignment tasks to enhance the policy models (Schulman et al., 2017), although it is computationally intensive. Consequently, several alternative, more efficient algorithms have been proposed, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024), Group Relative Policy Optimization Shao et al. (2024), and Rejection Sampling Fine-tuning (RAFT) (Dong et al., 2023). These approaches have also been widely adopted in state-of-the-art models like Llama-3 (Grattafiori et al., 2024), and Qwen-2 (Yang et al., 2024a).
@@ 2.2 Large Language Models and Retrieval Augmented Generation
In this project, we are the first to systematically construct RAG-scenario preference datasets and develop reward models, paving the way for evaluating and enhancing the generation quality of LLMs within the RAG framework.
Figure 2: An illustration of our data annotation method. Given a sample and two responses, we prompt o3-mini to provide a judgment based on each metric separately. We then aggregate the results and construct the pairs.
@@ 3 Dataset Construction
We construct our dataset based on existing RAG datasets to ensure its relevance and applicability. To reflect the diverse use cases of RAG scenarios, we include three common types: Question Answering, Data-to-Text and Summarization. Specifically, we use WebGLM (Liu et al., 2023), Yelp (Yelp, 2021), and XSum (Narayan et al., 2018) as experimental datasets, each dataset corresponding to one of the three RAG scenarios.
For the WebGLM dataset, LLMs are tasked with reasoning over web-retrieved reference data to answer real-world questions, generating concise responses in a few sentences. For the Yelp dataset, our experiments focus on data from the restaurant category, represented in JSON format. Each sample includes information such as a restaurant’s location and ambiance. Based on the structured JSON input, LLMs generate descriptive text about the restaurant. The XSum dataset contains diverse articles from the British Broadcasting Corporation (BBC), with models tasked with summarizing these articles. These three datasets cover a broad range of circumstances, ensuring that the reward model trained on them can significantly improve the development and evaluation of RAG systems. Table 1 presents the number of data samples used in our experiments. And examples of these data sets are presented in Table 2.
@@
While the response should be detailed and comprehensive, it should also be concise, relevant, and straight to the point. Striking the right balance between detail and brevity is essential to providing informative answers without overwhelming the user.
@@ 3.2 Dataset Labeling
We use o3-mini (OpenAI, 2025) to label the data. An illustration of our labeling methods is shown in Figure 2. Given a question and a pair of responses from different models, we prompt o3-mini to compare and select the preferred response.
Specifically, we ask o3-mini to compare the responses based on the four metrics outlined earlier, assessing them individually. In the prompt, we explicitly ask o3-mini to put heavier weights on hallucination and comprehensiveness metrics, as they are crucial to the answer quality, while the other two mainly improves the readability. After the o3-mini has made the individual judgments on the 4 metrics, it will generate an overall preference for the pair data based on the judgments above. We may end up getting a preferred answer with no hallucination but a bit verbose. And this approach acknowledges real-world scenarios where responses are rarely perfect, and trade-offs are often necessary.
@@ 4.1 Self Evaluation
Consist. 97.9 98.8 95.2 97.3
Table 3: The consistency rate for the self-evaluation of the constructed dataset. We measure whether the o3-mini agrees with the comparison results previously made.
Figure 3: An illustration of our self-evaluation method. For a sample of the constructed data pairs, we provide o3-mini with the previous judgment, and ask it to re-evaluate the comparison result.
In this subsection, we evaluate the consistency of the evaluations provided by o3-mini, a key metric commonly used to demonstrate the reliability of the annotations. To objectively assess the quality of our constructed dataset, we design a self-evaluation method that measures the consistency of o3-mini’s responses. An illustration of this method is shown in Figure 3. Specifically, for each prompt, a chosen response, and a rejected response in the dataset, we prompt o3-mini to revisit its previous comparison result and verify whether it maintains its original judgment. We define the consistency rate as the proportion of samples where the evaluation results remain unchanged across both assessments. In this experiment, we randomly select 1,000 samples for re-evaluation.
The results of this experiment are presented in Table 3. We observe a very high consistency rate across the three tasks and the overall consistency rate exceeds 97%. The results demonstrate that o3-mini produces stable and consistent labels for most of the data according to well-defined criteria, and reflects the quality of the dataset.
WebGLM Yelp XSum Avg
Consist. 0.79 0.80 0.83 0.81
Table 4: The consistency rate between human evaluation and the o3-mini labeled dataset.
@@ 4.2 Human Evaluation
UltraRM-13b (Cui et al., 2024) 71.0 77.4 79.2 75.9
llama-3-tulu-2-8b-uf-mean-rm (Ivison et al., 2024) 71.0 76.2 78.8 75.3
internlm2-7b-reward (Cai et al., 2024) 72.0 73.0 80.4 75.1
Eurus-RM-7b (Yuan et al., 2024) 71.8 74.4 77.8 74.7
FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2024) 71.6 72.4 77.8 73.9
Llama-3-OffsetBias-RM-8B (Park et al., 2024) 70.0 70.4 76.6 72.3
URM-LLaMa-3.1-8B (Lou et al., 2024) 66.8 70.2 76.0 71.0
QRM-Llama3.1-8B-v2 (Dorka, 2024) 68.4 68.0 74.6 70.3
GRM-Llama3.2-3B-rewardmodel-ft (Yang et al., 2024b) 64.6 71.6 73.4 69.9
Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) 64.8 68.6 72.8 68.7
Table 5: The evaluation results of the existing reward models on the 3 tasks. They achieve SOTA performance on chatting, safety, and reasoning evaluation, but do not excel in RAG tasks.
We also performed human evaluations to assess the alignment of AI annotations with human preferences. Specifically, we randomly select 100 samples with paired responses from each dataset and ask the annotators to evaluate using the same pipeline described in Section 3.2 and illustrated in Figure 2. Annotators compare the responses based on each metric and determine the preference pair. We calculate the agreement ratio between the human annotators and o3-mini in the preference pairs, and the results are shown in Table 4. We observe an overall agreement rate of 81%, with consistent agreement across the three tasks. This metric is comparable to the figures reported in Jin et al. (2024), where RAG is evaluated based on helpfulness and harmlessness criteria. These results highlight proprietary LLMs’ ability to effectively capture human preferences in assessing RAG response quality.